A ray of light impinges from air onto a block of ice (n = 1.309) at a 50.0angle of incidence. Assuming that this angle remains the same, find the radio of the angles of refraction (Ѳ2, ice = Ѳ2, water) when the ice turns to water (n = 1.333).


1
Expert's answer
2021-11-14T17:15:32-0500

For ice the angle of refraction is

"\theta_{ri}=\arcsin\frac{\sin\theta}{n_i}=\arcsin\frac{\sin50°}{1.309},\\\space\\
\theta_{ri}=36°."

For the molten ice:


"\theta_{rw}=\arcsin\frac{\sin\theta}{n_w}=\arcsin\frac{\sin50°}{1.333},\\\space\\
\theta_{rw}=35°."

The ratio:


"\frac{\theta_{ri}}{\theta_{rw}}=\frac{36°}{35°}=1.03."